OpenStudy (anonymous):

A rock is thrown in the air at time t=0. It's height h(t) in feet above the ground at time t in seconds is given by h(t)=-4t^2+24t+5. At what time (s) is the height of the rock 5 feet above the ground?

OpenStudy (saifoo.khan):

h(t)=-4t^2+24t+5 Now since the height is 5, 5 = -4t^2+24t+5 Solve for "t".

OpenStudy (saifoo.khan):

You''be left over with: -4t^2+24t = 0 Now solve for t.

OpenStudy (anonymous):

it means tht...... 5=-4t^2+24t+5 Subtrating by 5 on both sides, we get -4t^2+24t=0 4t is common here so by factorizing we get -4t(t-6)=0 therefore -4t=0 or t-6=0 t=0(rejected) or t=6 the time taken =6

OpenStudy (anonymous):

Thank you, I thought I was going crazy with this problem. It's great to know that I was not too far off from the right path. Thank you, you all are life savers!!!!!!!

OpenStudy (anonymous):

Its ok we r here to help you!!!!!